A hospital director is told that 79% of the treated patients are insured. The director wants to test the claim that the percentage of insured patients is less than the expected percentage. A sample of 240 patients found that 180 were insured. Find the value of the test statistic. Round your answer to two decimal places.
Answer #1

Given that, n = 240 and x = 180

sample proportion (hat p) = x/n = 180/240 = 0.75

The null and the alternative hypotheses are,

Ho : p 0.79

Ha p<0.79

Test statistic is,

0.75-0.79__15214 p-P 1-2) 0.79 (1-0.79

Test statistic is, z = -1.52